# Causal Inference in Energy Demand Prediction

### v1

@@ 2.3. Humidity
Figure 4. The Distribution of Humidity Across Local Hours
@@ 3.1. Model Setup
$+\underbrace{\sum_{j=1}^{n}(d_{j}\sin(2\pi jH/24)+d_{j}^{*}\cos(2\pi jH/24))}_{\text{daily temperature cycle}}+T_{\text{base}}+\underbrace{e_{\text{temp}}}_{\text{noise}},$
where $T_{\text{base}}$ is a constant baseline and $e_{\text{temp}}\sim\mathcal{N}(0,\sigma^{2})$ is Gaussian random noise.
Humidity, solar radiation and wind speed are sampled from independent Gaussian distributions:
(7) $\text{Humidity}\sim\mathcal{N}(\mu_{\text{humid}},\sigma_{\text{humid}}^{2}),$
$\text{Radiation}\sim\mathcal{N}(\mu_{\text{rad}},\sigma_{\text{wind}}^{2}),$
$\text{Wind}\sim\mathcal{N}(\mu_{\text{wind}},\sigma_{\text{wind}}^{2}).$

### v2

@@ 2.3. Humidity
(a) January
(b) February
Figure 4. The Distribution of Humidity Against Month and Hour
@@ 3.1. Model Setup
$+\underbrace{\sum_{j=1}^{n}(d_{j}\sin(2\pi jH/24)+d_{j}^{*}\cos(2\pi jH/24))}_{\text{daily temperature cycle}}$ $+w\cdot\text{Rad}+T_{\text{base}}+\underbrace{e_{\text{temp}}}_{\text{noise}},$
where Rad refers to solar radiation, $T_{\text{base}}$ is a constant baseline and $e_{\text{temp}}\sim\mathcal{N}(0,\sigma^{2})$ is Gaussian random noise.
We sample humidity from a beta distribution $Beta(\alpha,\beta)$ whose parameters are determined by month and hour.
(7) $\alpha=\theta\sin(\frac{\pi H}{12})+\theta^{*}\cos(\frac{\pi H}{12})+\alpha_{0}$
$\beta=\phi\sin(\frac{\pi M}{6})+\phi^{*}\cos(\frac{\pi M}{6})+\beta_{0}-\alpha$
The choice of distribution is based on empirical study, which reveals that the concentration and skewness of the humidity distribution are influenced by month and hour, see figure 4, for example, where the shape of the humidity distribution varies greatly by month and hour.
Solar radiation is sampled from a normal distribution and is cast to non-negative values afterwards. The mean of the normal distribution is defined as a function of month and hour, where the effects of month and hour are factorized into separate periodic components.
(8) $\mu_{\text{Rad}}=\left\{\begin{aligned} &(p\sin(\frac{\pi M}{12})+q)\sin(\pi\frac{H-H_{\text{sunrise}}(M)}{H_{\text{sunset}}(M)-H_{\text{sunrise}}(M)}),\\ &\hskip 39.83368ptif\ H_{\text{sunrise}}(M)\leq H\leq H_{\text{sunset}}(M),\\ &0,\hskip 28.45274ptotherwise.\end{aligned}\right$.
where $H_{\text{sunrise}}(M)$ and $H_{\text{sunset}}(M)$ are respectively the local sunrise and sunset time, averaged over the month $M$. The form of equation (8) is carefully selected based on the empirical results, as shown in figure 12.
(a) Radiation vs. Month
(b) Radiation vs. Hour
Figure 12. Solar Radiation vs. Calendar Variables
Wind speed is sampled directly from normal distribution $\mathcal{N}(\mu_{wind},\sigma_{wind})$ and then cast to non-negative values.